# Reformulating Conversational Recommender Systems as Tri-Phase Offline Policy Learning

### v1

@@ 5.3. Performance Comparison (RQ1)
  • The Tri-Phase Offline Policy Learning-based Conversational Recommender System (TPCRS) framework demonstrates superior recommendation performance across all datasets compared to the item-centric user simulator-based baselines (EAR, SCPR, UNICORN, MCMIPL, AVPPL) as well as the simulator-free baselines (Abs Greedy and Max Entropy).
  • In smaller datasets like LastFM, Abs Greedy’s exhibits competitive performance, showing the potency of Preference Estimation User Model (PEUM) component within the TPCRS framework. However, it struggles in the Yelp and Amazon-Book datasets, which contain a large number of items. In comparison, Max Entropy’s poor performance across all datasets highlights the importance of policy learning to capture personalized and dynamic user preferences.
  • AVPPL and TPCRS outperform other CRS models as they conduct meticulous user behavior assumptions tailored to policy learning, better capturing realistic user preference modeling.
  • In Yelp, TPCRS leverages long-term dynamic preferences more effectively despite weaker short-term performance. This shows TPCRS’s strength in capturing evolving user preferences long-term. The abundance of items per attribute in the Yelp dataset leads to the TPCRS seeking asking more rounds of information before providing precise recommendations.
5.4. Evaluating Key Design in TPCRS (RQ2)
Table 3. Evaluation of the PEUM component and different strategies in the TPCRS framework.
@@ 5.4. Evaluating Key Design in TPCRS (RQ2)
(a) - w/o PEUM Item 0.682 8.15 0.257 12.43 0.349 11.37
(b) - w/o PEUM Attr. 0.598 8.89 0.224 13.12 0.360 11.27
(c) - w/o PEUM Both 0.690 8.23 0.313 12.32 0.393 11.03
@@
5.4.1. Ablation Study on PEUM
The proposed Preference Estimation User Model (PEUM) module, a key TPCRS framework component, captures dynamic user preferences by jointly modeling items and attributes. We compare the full PEUM model with two graph-based variants (Deng et al., 2021; Zhang et al., 2022, 2023): 1) without the Item Preference Ranking module, and 2) without the Attribute Preference Ranking module.
Table 3 (a-c) shows that removing either the item or attribute preference prediction leads to performance degradation across the Yelp and Amazon-Books datasets. This highlights the importance of PEUM’s joint modeling in constructing an accurate and adaptive preference action space for the downstream RL-based policy. Interestingly, the graph-based preference modeling approach (row c) performs reasonably well on the LastFM dataset. This is likely due to the denser attribute information available in LastFM, which plays to the strengths of the graph-based method. However, our full PEUM model still achieves superior average turn counts, underscoring its efficiency in streamlining recommendations compared to the graph-based baseline.
@@ 5.4.2. Sampling Strategies of Controllable User Simulation
Another key TPCRS component, the Controllable User Simulation module, provides a realistic evaluation by modeling the personalized and evolving nature of user preferences. A crucial aspect of this module is the sampling strategies used to generate user feedback during the conversational recommendation evaluation.
@@ 5.5. Adaptability Evaluation (RQ3)
Figure 4. Evaluating TPCRS policies under varying personalized preference ($\alpha$) and preference evolution rate ($\Delta\lambda$) in the Controllable User Simulation (RQ3).
A key contribution of the TPCRS framework is its ability to adapt to diverse and dynamically evolving user preferences. To assess this adaptability, we leverage the Controllable User Simulation module to create varied user preference environments.
@@
Increasing $\alpha$ emphasizes personalized preferences over dynamic adaptation to target item. The results show that higher $\alpha$ leads to performance declines on Yelp and Books, as policies become overly reliant on static preferences. However, LastFM exhibits stable performance, suggesting personalized preferences are sufficient for accurate recommendations in this domain. Elevating $\Delta\lambda$ introduces a more dynamic preference landscape, testing the policies’ ability to adapt. The success rate remains largely stable, underscoring the TPCRS framework’s robustness in reconciling evolving preferences with recommendation convergence.
These results showcase the TPCRS framework’s effectiveness in aligning with diverse user preference characteristics, transcending the limitations of traditional item-centric simulators.
@@ 5.6. Exploring User Preference Evolution (RQ4)
A key advantage of the TPCRS framework is its ability to capture the dynamic and personalized nature of user preferences, in contrast to the static, item-centric assumptions of traditional CRS simulators. To validate this capability, we examine how the alignment between the user’s evolving preferences and the attributes of the target item changes over the course of the conversational interaction.
@@
The results show that while the initial match rates differ across datasets, the overlap consistently improves as the conversation progresses. This pattern holds true regardless of the level of personalization (controlled by the $\alpha$ parameter), highlighting the TPCRS framework’s adaptive capacity in modeling evolving user preferences. As the system inquires about user preferences and makes recommendations, it is able to successfully align the user’s interests with the target item attributes.
@@ 6. Conclusion
The Tri-Phase Offline Policy Learning-based Conversational Recommender System (TPCRS) framework represents a significant advancement in conversational recommender systems. by addressing the critical limitations of existing approaches. By decoupling the training and evaluation phases, TPCRS introduces an effective offline policy learning method that effectively mitigates the overfitting issue associated with traditional simulator-based approaches.
The explicit modeling of user preference dynamics and personalization within the Preference Estimation User Model empowers the CRS policies to better capture and adapt to real-world user behaviors. The introduction of the Controllable User Simulation module provides a more realistic and comprehensive evaluation environment, enabling a thorough assessment of policy performance across diverse user scenarios.
The key innovations of TPCRS, including the tri-phase architecture, the multi-task user preference modeling, and the user-centric simulation, pave the way for more effective and robust conversational recommender systems. These advancements allow CRS policies to better understand and cater to individual user needs, leading to enhanced user experiences and increased adoption of CRS in real-world applications.
@@ 7. Acknowledgements
This work is supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20241643, and the Fundamental Research Funds for the Central Universities of China (Grant No.PA2024GDSK0107). This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.

### v2

@@ 5.3. Performance Comparison (RQ1)
  • The Tri-Phase Offline Policy Learning-based Conversational Recommender System (TCRS) framework demonstrates superior recommendation performance across all datasets compared to the item-centric user simulator-based baselines (EAR, SCPR, UNICORN, MCMIPL, VPPL) as well as the simulator-free baselines (Abs Greedy and Max Entropy).
  • In smaller datasets like LastFM, Abs Greedy’s exhibits competitive performance, showing the potency of Conversational User Model (CUM) component within the TCRS framework. However, it struggles in the Yelp and Amazon-Book datasets, which contain a large number of items. In comparison, Max Entropy’s poor performance across all datasets highlights the importance of policy learning to capture personalized and dynamic user preferences.
  • VPPL and TCRS outperform other CRS models as they conduct meticulous user behavior assumptions tailored to policy learning, better capturing realistic user preference modeling.
  • In Yelp, TCRS leverages long-term dynamic preferences more effectively despite weaker short-term performance. This shows TCRS’s strength in capturing evolving user preferences long-term. The abundance of items per attribute in the Yelp dataset leads to the TCRS seeking asking more rounds of information before providing precise recommendations.
5.4. Evaluating Key Design in TCRS (RQ2)
Table 3. Evaluation of the CUM component and different sampling strategies in the TCRS framework.
@@ 5.4. Evaluating Key Design in TCRS (RQ2)
(a) - w/o CUM Item 0.682 8.15 0.257 12.43 0.349 11.37
(b) - w/o CUM Attr. 0.598 8.89 0.224 13.12 0.360 11.27
(c) - w/o CUM Both 0.690 8.23 0.313 12.32 0.393 11.03
@@
5.4.1. Ablation Study on CUM
The proposed Conversational User Model (CUM) module, a key TCRS framework component, captures dynamic user preferences by jointly modeling items and attributes. We compare the full CUM model with two graph-based variants (Deng et al., 2021; Zhang et al., 2022, 2023): 1) without the Item Preference Ranking module, and 2) without the Attribute Preference Ranking module.
Table 3 (a-c) shows that removing either the item or attribute preference prediction leads to performance degradation across the Yelp and Amazon-Books datasets. This highlights the importance of CUM’s joint modeling in constructing an accurate and adaptive preference action space for the downstream RL-based policy. Interestingly, the graph-based preference modeling approach (row c) performs reasonably well on the LastFM dataset. This is likely due to the denser attribute information available in LastFM, which plays to the strengths of the graph-based method. However, our full CUM model still achieves superior average turn counts, underscoring its efficiency in streamlining recommendations compared to the graph-based baseline.
@@ 5.4.2. Sampling Strategies of Controllable User Simulation
Another key TCRS component, the Controllable User Simulation module, provides a realistic evaluation by modeling the personalized and evolving nature of user preferences. A crucial aspect of this module is the sampling strategies used to generate user feedback during the conversational recommendation evaluation.
@@ 5.5. Adaptability Evaluation (RQ3)
Figure 4. Evaluating TCRS policies under varying personalized preference ($\alpha$) and preference evolution rate ($\Delta\lambda$) in the Controllable User Simulation (RQ3).
A key contribution of the TCRS framework is its ability to adapt to diverse and dynamically evolving user preferences. To assess this adaptability, we leverage the Controllable User Simulation module to create varied user preference environments.
@@
Increasing $\alpha$ emphasizes personalized preferences over dynamic adaptation to target item. The results show that higher $\alpha$ leads to performance declines on Yelp and Books, as policies become overly reliant on static preferences. However, LastFM exhibits stable performance, suggesting personalized preferences are sufficient for accurate recommendations in this domain. Elevating $\Delta\lambda$ introduces a more dynamic preference landscape, testing the policies’ ability to adapt. The success rate remains largely stable, underscoring the TCRS framework’s robustness in reconciling evolving preferences with recommendation convergence.
These results showcase the TCRS framework’s effectiveness in aligning with diverse user preference characteristics, transcending the limitations of traditional item-centric simulators.
@@ 5.6. Exploring User Preference Evolution (RQ4)
A key advantage of the TCRS framework is its ability to capture the dynamic and personalized nature of user preferences, in contrast to the static, item-centric assumptions of traditional CRS simulators. To validate this capability, we examine how the alignment between the user’s evolving preferences and the attributes of the target item changes over the course of the conversational interaction.
@@
The results show that while the initial match rates differ across datasets, the overlap consistently improves as the conversation progresses. This pattern holds true regardless of the level of personalization (controlled by the $\alpha$ parameter), highlighting the TCRS framework’s adaptive capacity in modeling evolving user preferences. As the system inquires about user preferences and makes recommendations, it is able to successfully align the user’s interests with the target item attributes.
@@ 6. Conclusion
The Tri-Phase Offline Policy Learning-based Conversational Recommender System (TCRS) framework represents a significant advancement in conversational recommender systems. by addressing the critical limitations of existing approaches. By decoupling the training and evaluation phases, TCRS introduces an effective offline policy learning method that effectively mitigates the overfitting issue associated with traditional simulator-based approaches.
The explicit modeling of user preference dynamics and personalization within the Conversational User Model empowers the CRS policies to better capture and adapt to real-world user behaviors. The introduction of the Controllable User Simulation module provides a more realistic and comprehensive evaluation environment, enabling a thorough assessment of policy performance across diverse user scenarios.
The key innovations of TCRS, including the tri-phase architecture, the multi-task user preference modeling, and the user-centric simulation, pave the way for more effective and robust conversational recommender systems. These advancements allow CRS policies to better understand and cater to individual user needs, leading to enhanced user experiences and increased adoption of CRS in real-world applications.
@@ 7. Acknowledgements
This work is supported by the Fundamental Research Funds for the Central Universities of China (WK2100000053, PA2024GDSK0107), the National Natural Science Foundation of China (No. 62402470), and the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20241643. This research is also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.